# Early Jamming Detection in Mobile Indoor Scenarios via Deep Learning

### v1

@@ Abstract
The current state of the art on jamming detection relies on link-layer metrics. A few examples are the bit-error rate, the packet delivery ratio, the throughput, and the increase of the signal-to-noise ratio. As a result, these techniques can only detect jamming ex-post, i.e., once the attack has already taken down the communication link. These solutions are unfit in mobile scenarios, e.g., drones, which might lose the link to the remote controller, being unable to predict the attack.
Our solution is rooted in the idea that a drone flying against a jammed area is experiencing an increasing effect of the jamming. Therefore, drones might use this phenomenon to detect jamming early, i.e., before it completely disrupts the communication link. Such an approach would allow drones and possibly their pilots to make an informed decision and maintain full control of the navigation, providing security and safety.
In this paper, we propose Bloodhound$+$, a solution for early jamming detection on mobile devices. Our approach analyzes raw physical-layer information (I-Q samples) acquired from the channel. We assemble this information into grayscale images, and we use sparse autoencoders to detect image anomalies caused by jamming attacks. To test our solution against a wide set of configurations, we acquired a large dataset of indoor measurements using multiple hardware, jamming strategies, and communication parameters. Our results indicate that Bloodhound$+$ can detect indoor jamming up to 20 meters away from the jamming source at the minimum available relative jamming power, with a minimum accuracy of 99.7%. Our solution is also robust to various sampling rates adopted by the jammer, as well as to the type of signal used for jamming.
@@
Wireless Security; Artificial Intelligence for Security; Drones Security.
@@ I Introduction
The adoption of remotely piloted and (semi) autonomous robots and drones in everyday business is increasing at a high pace. Today, such technologies are increasingly used both outdoors, e.g., for disaster management, search and rescue, and goods delivery [1], and indoors, e.g., for inventory management, intra-logistics of items, inspection and surveillance [2]. In this context, leading market analysis companies estimate a current market value of up to 100.37 billion USD by 2029, with a compound annual growth rate of 25.5% in the next years [3].
@@
As their role becomes more central, drones are increasingly the main target of several cybersecurity attacks. In particular, due to the reliance on wireless channels for communication, video streaming, and telemetry, the easiest way attackers can disrupt drones’ operation is through jamming attacks [4]. Jamming can significantly disrupt wireless communications in a given area through injection of high-power noise into the same channel used by legitimate communication parties [5]. In such cases, depending on the firmware onboard, the drone might return to the starting point of the mission, land, or even crash, with potential hazards for people in the area, especially for indoor applications [6].
@@
At the time of writing, the most available solutions for jamming detection work by identifying sudden drops in the Bit-Error Rate (BER), the Packet Delivery Rate (PDR), or the Signal to Noise Ratio (SNR) of the communication link. Some of them also use the spectrogram of the signal in the Physical (PHY) layer, looking for sudden anomalies (see Section VI for an overview). Such solutions work reliably and effectively when applied to static deployments. However, they can detect jamming mainly ex-post, i.e., once jamming has already disrupted the regular operations of the communication link. The application of such approaches on drones would trigger the default actions listed above, possibly causing hazards to the drone and the surrounding people. In this context, mobile devices such as drones typically experience an increasing effect of the jamming (high BER, low PDR, low SNR) while approaching the jammed area. Drones might exploit such phenomena to deploy a solution able to detect jamming earlier, i.e., before entering an area where the high-power noise injected by the jammer completely disrupts the communication link. Such a solution is especially critical for indoor applications, where people and drones often work together in limited space.
@@
Contribution. In this paper, we propose BloodHound$+$, an innovative solution for early jamming detection in mobile indoor scenarios leveraging state-of-the-art Deep Learning (DL) techniques. BloodHound$+$ allows to carry out jamming detection by converting raw PHY data, that is, I-Q samples, into images while detecting anomalies in their shape by resorting to autoencoders.
@@
When applied to mobile entities, for example, autonomous or remotely piloted vehicles and drones, BloodHound$+$ can detect the approach of a jammed area well before jamming could completely disrupt the communication link, allowing the remote entity to maintain full control. To verify the effectiveness of our proposed approach, we carried out an extensive measurement campaign, involving multiple hardware devices (multiple Ettus Research X310 and LimeSDR radios), communication link configurations, and jamming conditions. Using such measurements, we tested the effectiveness of BloodHound$+$ and other competing approaches for the early detection of jamming. Our results show that BloodHound$+$ can detect jamming well earlier compared to benchmark solutions, for example, with an accuracy of $0.997$ when the adversary jams at a distance of $10$ m from the target with a Relative Jamming Power (RJP) of $0.1$. Our solution is also very robust to: (i) the distance from the jammer, (ii) the training set size, (iii) the number of acquired samples, (iv) the sampling ratios at the jammer and the receiver, (v) the type of jamming signal (tone, Gaussian, or deceptive), as well as, (vi) the adoption of different jamming hardware and radio types.
@@
We note that this contribution extends and completes our previous work published in [7], by providing the following new content.
@@
  • We consider a stronger adversary model, assuming that the adversary is aware of the sampling rate and the modulation techniques used by the legitimate communication link. Such knowledge allows the adversary to optimize the parameters of the jamming attack, in order to boost its effectiveness and avoid detection at the same time.
@@
  • We design a new optimized methodology for jamming detection, based on a one-class classifier of black and white images extracted from raw I-Q samples using sparse autoencoders.
@@
  • We experimentally compare our new methodology with the ones proposed in [7] and [8], showing remarkable performances and improvements with reference to the RJP at the receiver, distance from the jammer, number of I-Q samples per image, training set size, and invariance to the hardware used for training.
@@
  • We provide additional results on a new dataset, gathered using new hardware, namely, the LimeSDR.
@@
  • We provide new results obtained on the data gathered by using various sampling rates at the jammer and receiver.
@@
  • We provide new results on the usage of a new jamming strategy, i.e., deceptive jamming using the same modulation ( Binary Phase-Shift Keying (BPSK)) and signal of the legitimate communication link.
@@
Roadmap. The rest of this paper is organized as follows. Section II introduces preliminary notions; Section III describes the scenario and adversarial model; Section IV provides the rationale and details of BloodHound$+$; Section V discusses our extensive measurement campaign and performance assessment of our solution, and finally Section VII draws the conclusion and outlines future work.
II Preliminaries
@@
In this section, we introduce preliminary notions that are useful to the readers of this manuscript, i.e., digital modulation techniques (Section II-A) and autoencoders (Section II-B).
II-A Digital Modulation
@@
Digital modulation schemes adopted in wireless communication systems preprocess baseband signals to make them suitable for transmission at high frequencies [9]. Typically, modulation techniques divide the bitstream to be transmitted into two orthogonal components, namely the I vector and the Q vector, linked in a complex value of type $I+jQ$, where the I vector is the real component and the Q vector is the imaginary component. Due to their orthogonality, such components can be transmitted together on the wireless channel without interfering and can also be recovered and assembled at the receiver to reconstruct the original bit-stream. In this context, a usual way to represent complex I-Q signals is through the I-Q plane, as shown in Fig. 1.
@@
In particular, the number of expected I-Q values at the receiver (denoted as $n$) indicates the amount of bits that can be recovered through a single complex I-Q value. In general, we can recover $\log_{2}{n}$ bits from $n$ symbols, and thus, with reference to Fig. 1 showing the I-Q plane of a BPSK modulation, we can recover $n=2$ bits, i.e., $\left[i=-1,q=0\right]$ (b=0) and $\left[i=1,q=0\right]$ (b=1). We also notice that the values of the I-Q samples at the receiver always differ from those at the transmitter because of noise introduced by the hardware components of the devices and the fluctuations of the wireless channel between the transmitter and the receiver. To recover the original transmitted symbol, the receiver associates to the received I-Q sample the expected symbol whose distance from the received one is the shortest. Thus, the higher the impact of noise, the higher the chance that a received I-Q sample is associated with the wrong expected symbol, leading to an error and therefore to a higher BER. At the same time, with a given noise profile affecting the communication channel, the higher the modulation order $n$, the higher the amount of expected symbols and, thus, the higher the BER. This is the rationale for the use of lower-order modulation schemes for communication links affected by high levels of noise, such as satellite transmissions and mobile indoor applications. The intuition driving our work is that the collective displacement of I-Q samples from the expected one can be used to discriminate the presence of various levels of intentional interference, i.e., jamming, affecting the communication link. We will provide more details about our approach in Section IV.
II-B Autoencoders
@@
where the symbol $\circ$” represents the composition operator, i.e., $B\circ A(\mathbf{x})=B(A(\mathbf{x}))$, $E$ represents the expectation of the distribution of the input $\mathbf{x}$, ’$A(\mathbf{x})$’ the encoded version of the input, known as the bottleneck of the autoencoder when $p<d$, and finally $\Delta$ is the reconstruction loss function, which measures the distance between the input of the ANN and the reconstruction of the input [10]. For our purposes, in line with many scientific contributions such as [10] and [11], $\Delta$ is the mean-squared-error (mse) function, as defined in Eq. 2 on two reference distributions $x$ and $y$.
@@
Traditionally, autoencoders have been used mostly for image generation and, in particular, for creating sets of images similar to the input ones. However, in the cybersecurity research domain, they are mainly used for anomaly detection. Let $c$ be an autoencoder trained on samples taken from a probability distribution $P$. Next, let $Q$ be a probability distribution such that $P\neq Q$. Then, we expect $c$ to have a smaller reconstruction error when tested on unseen samples from $P$ than when tested on unseen samples from $Q$. Therefore, the magnitude of the reconstruction error of $c$ on an unseen sample is a measure of the probability that such an unseen sample is not sampled from $P$, but from another distribution [11]. Consequently, we can define a specific error value $\tau$ as a classification boundary. All samples in which the auto-encoder $c$ makes a higher error than $\tau$ can be classified as being ’not from $P$’. In the literature, $\tau$ is often referred to as a threshold [12, 13].
@@
In this work, we use autoencoders to build a statistical profile of the channel experienced between the mobile transmitter and the receiver under regular operating conditions. We provide more details in Section IV.
III System and Adversary Model

### v4

# Jamming Detection in Low-BER Mobile Indoor Scenarios via Deep Learning

@@ Abstract
The current state of the art on jamming detection relies on link-layer metrics. A few examples are the bit-error rate (BER), the packet delivery ratio, the throughput, and the signal-to-noise ratio (SNR). As a result, these techniques can only detect jamming ex-post, i.e., once the attack has already taken down the communication link. These solutions are unfit for mobile devices, e.g., drones, which might lose the connection to the remote controller, being unable to predict the attack. Our solution is rooted in the idea that a drone unknowingly flying toward a jammed area is experiencing an increasing effect of the jamming, e.g., in terms of BER and SNR. Therefore, drones might use the abovementioned phenomenon to detect jamming before the increase of the BER and the decrease of the SNR completely disrupt the communication link. Such an approach would allow drones and their pilots to make informed decisions and maintain complete control of navigation, enhancing security and safety. This paper proposes Bloodhound$+$, a solution for jamming detection on mobile devices in low-BER regimes. Our approach analyzes raw physical-layer information (I-Q samples) acquired from the wireless channel. We assemble this information into grayscale images and use sparse autoencoders to detect image anomalies caused by jamming attacks. To test our solution against a broad set of configurations, we acquired a large dataset of indoor measurements using multiple hardware, jamming strategies, and communication parameters. Our results indicate that Bloodhound$+$ can detect indoor jamming up to 20 meters from the jamming source at the minimum available relative jamming power, with a minimum accuracy of 99.7%. Our solution is also robust to various sampling rates adopted by the jammer and to the type of signal used for jamming.
@@
Wireless Security; Artificial Intelligence for Security; Drones Security; Mobile Security.
@@ I Introduction
Drones, a.k.a. Unmanned Aerial Vehicles, represent an evolution of the Internet of Things (IoT) paradigm. Since we define the IoT as a network of physical objects (things) that are embedded with sensors, software, and other technologies to connect and exchange data with other devices and systems over the Internet [1],[2], it is immediate to include drones in the IoT domain. They are battery-powered devices, featuring limited onboard storage space and integrating Central Processing Units with heterogeneous computational capabilities. Recent papers even coined the notion of Flying IoT to specifically identify drones as a specification of the IoT paradigm [3][4][5][6].
@@
Today, drones are increasingly used both outdoors, e.g., for disaster management, search and rescue, and goods delivery [7], and indoors, e.g., for inventory management, intra-logistics of items, inspection and surveillance [8], and leading companies such as IKEA and Amazon are actively experimenting products for warehouse management and home surveillance [9][10]. In this context, leading market analysis companies estimate a current market value of up to 100.37 billion USD by 2029, with a compound annual growth rate of 25.5% in the next years [11].
@@
As their role becomes more central, drones are increasingly the main target of several cybersecurity attacks. In particular, due to the reliance on wireless channels for communication, video streaming, and telemetry, attackers can easily disrupt drones’ operation through jamming attacks [12]. Jamming can significantly disrupt wireless communications in a given area by injecting high-power noise into the same channel used by legitimate communication parties [13]. Depending on the firmware onboard, the drone might return to the mission’s starting point, land, or even crash, with potential hazards for people in the area, especially for indoor applications [14].
@@
At the time of writing, most of the available solutions for jamming detection work by identifying the deterioration of the Bit-Error-Rate (BER), the Packet Delivery Rate (PDR), or the Signal to Noise Ratio (SNR) of the communication link. Some of them also use the spectrogram of the signal in the Physical (PHY) layer, looking for sudden anomalies (see Section II for an overview). Such solutions work reliably and effectively when applied to static deployments. However, they can detect jamming mainly ex-post, i.e., once jamming has already disrupted the regular operations of the communication link. Applying such approaches to drones would trigger the default actions listed above, possibly causing hazards to the drone and surrounding people. In this context, mobile devices such as drones typically experience an increasing effect of jamming (increasingly high BER, low PDR, low SNR) while approaching the jammed area. Drones might exploit the mentioned phenomena to deploy a solution to detect jamming in a low-BER regime, i.e., before entering an area where the high-power noise injected by the jammer completely disrupts the communication link. Such a solution is especially critical for indoor applications, where people and drones often work together in limited space.
@@
Contribution. In this paper, we propose BloodHound$+$, an innovative solution for jamming detection in low-BER mobile indoor scenarios leveraging state-of-the-art Deep Learning (DL) techniques. BloodHound$+$ allows to carry out jamming detection by converting raw PHY data, that is, I-Q samples, into images while detecting anomalies in their shape by resorting to autoencoders.
@@
Although our approach can apply to any mobile device, we specifically consider the case of drones operating indoors, e.g., for warehouse inspections [15], inspired by the recent deployment of real-world use cases, such as the one managed by IKEA [9] and Amazon [10]. When applied to autonomous or remotely piloted vehicles and drones, BloodHound$+$ can detect the approach of a jammed area with very low BER values, allowing the remote entity to detect jamming while maintaining complete control of the communication link. To verify the effectiveness of our proposed approach, we conducted an extensive measurement campaign emulating UAVs through Ettus Research X310 and LimeSDR radios, using multiple hardware devices (multiple Ettus Research X310 and LimeSDR radios), communication link configurations, and jamming conditions. Using such measurements, we tested the effectiveness of BloodHound$+$ and other competing approaches for detecting jamming in a low-BER regime. Our results show that BloodHound$+$ can detect jamming in scenarios with a lower BER ($\approx 1e-6$) compared to benchmark solutions, e.g., with an accuracy of $0.997$ when the adversary jams at a distance of $10$ m from the target with a Relative Jamming Power (RJP) of $0.1$. Our solution is also very robust to: (i) the distance from the jammer, (ii) the training set size, (iii) the number of acquired samples, (iv) the sampling ratios at the jammer and the receiver, (v) the type of jamming signal (tone, Gaussian, or deceptive), as well as (vi) the adoption of different jamming hardware and radio types. We envision that BloodHound$+$ can be used during regular operations of the UAV by acquiring physical layer information from the (already existing) UAV communications channel. In addition, it can be activated when desired, so as not to cause significant overhead on the UAV.
@@
This contribution extends and completes our previous work published in [16] by providing the following new content.
@@
  • We consider a stronger adversary model, assuming that the adversary knows the sampling rate and modulation techniques of the legitimate communication link. This knowledge allows the adversary to optimize the parameters of the jamming attack to boost its effectiveness and avoid detection simultaneously.
@@
  • We design a new optimized methodology for jamming detection based on a one-class classifier of black and white images extracted from raw I-Q samples using sparse autoencoders.
@@
  • We experimentally compare our new methodology with those proposed in [16] and [17], showing remarkable performances and improvements concerning the RJP at the receiver, distance from the jammer, number of I-Q samples per image, training set size, and invariance to the hardware used for training.
@@
  • We provide additional results on a new dataset gathered using new hardware, namely, the LimeSDR.
@@
  • We provide new results on the data collected using various sampling rates at the jammer and receiver.
@@
  • We provide new results on using a new jamming strategy, i.e., deceptive jamming using the same modulation ( Binary Phase-Shift Keying (BPSK)) and signal of the legitimate communication link.
We acknowledge that BloodHound$+$ takes inspiration from anomaly detection strategies applied in other research domains, e.g., intrusion detection and computer vision [18]. However, to the best of our knowledge, none of the contributions in the current literature provided a structured methodology to apply such strategies to detect jamming attacks on the wireless RF spectrum. Also, as described in more detail in Sect. II, none of the contributions focused on low-BER regimes, investigating how to detect jamming while still maintaining remote communication capabilities.
@@
Roadmap. The rest of this paper is organized as follows. Section III introduces preliminary notions; Section IV describes the scenario and adversarial model; Section V provides the rationale and details of BloodHound$+$; Section VI discusses our extensive measurement campaign and performance assessment of our solution, and finally Section VII draws the conclusion and outlines future work.
II Related Work
[…]
Ref.
Jamming Detection
Metric
Jamming Detection
Technique
Robustness to
Jamming Distance
Robustness to
Jamming Signal Type
Jamming Detection
in Low-BER Regime
[19]
RSS
Geometric and Arithmetic Mean ratio
[…]
[20]
PDR
Query-based procedure
[…]
[21]
Carrier-to-Noise density power ratio
Sum-of-Squares Paradigm
[…]
[22]
Coherence blocks
Generalized Likelihood Ratio Test
[…]
[23]
RSS
Predetermined knowledge, Error Correcting Codes, and Limited node Wiring
[…]
[24], [25]
PDR
Code Tree
[…]
[26]
Re-transmissions
Statistics-based
[…]
[27]
Automatic Gain Control (AGC)
Static tests
[…]
[28]
PDR
Random forests
[…]
[29]
PDR
Channel probing
[…]
[30]
Retransmisions
Message Invalidation Ratio
[…]
[31]
OFDM parameters
Machine Learning (ML)
[…]
[32]
Power spectral density, spectrogram, raw constellation
ML
[…]
[33]
Spectrogram
ML
[…]
[34]
Nonlinear alternating current
CUSUM
[…]
[35]
PHY, Radio Link Control and Packet Data Convergence Control parameters
LSTM
[…]
[36]
I-Q
Autoencoders
[…]
[17]
I-Q
CNN
[…]
[16]
I-Q
CNN
[…]
BloodHound$+$
I-Q
Sparse Autoencoders
[…]
In the scientific community, jamming detection is usually achieved by applying various types of analysis on one or more metrics extracted from the primary communication link.
Regarding the metrics, several parameters have been analyzed, such as the Received Signal Strength (RSS) of the signals [19], the PDR as in [20] and [28], the Carrier-to-Noise density power ratio [21], retransmission attempts [26], and [29], the packet re-transmission profile, as in [30], or modulation-specific metrics, such as for Orthogonal Frequency Division Multiplexing (OFDM) in [31]. Such metrics have been used in several scenarios and communication technologies, e.g., Massive MIMO [22], Wireless Sensor Networks (WSN[23], GPS [27], IEEE 802.11 [28], and spread spectrum-based communication technologies [24][25].
At the same time, due to the increasing popularity of Artificial Intelligence (AI), ML and DL approaches have been recently used extensively for detecting ongoing jamming. Such tools include Convolutional s such as in [32] and [33], genetic algorithm-based Cumulative Sum (CUSUM) methods such as in [34], Bayesian networks such as in [35] and, finally, autoencoders such as in [36]. All such approaches utilize as the main source of information the PHY layer of the communication stack due to its direct relationship with the wireless channel, where jamming occurs. However, although some of the contributions cited analyzed the performance of the proposed jamming detection technique with low SNR, none of them considered the BER of the communication link. As a result, the proposed approaches mostly confirm that the root cause of the drop in BER is jamming. However, they cannot detect such attacks even when the jamming effect is so low as not to significantly affect the BER of the communication link. As explained above, such consideration is particularly relevant in mobile scenarios for remotely controlled equipment, not to lose control of the mobile entity completely before detecting jamming. In this context, the only contribution to the literature achieving such a property is our previous proposal in [16]. As shown in Section VI, the methodology shown in this paper significantly outperforms both the solution proposed in [16] and the improvements of such a methodology, as the one proposed in [17]. We summarize our comparison with the current literature in Tab. I.
III Preliminaries
@@
In this section, we introduce preliminary notions that are useful to the readers of this manuscript, i.e., digital modulation techniques (Section III-A) and autoencoders (Section III-B).
III-A Digital Modulation
@@
Digital modulation schemes adopted in wireless communication systems preprocess baseband signals to make them suitable for transmission at high frequencies [41]. Typically, modulation techniques divide the bit-stream to be transmitted into two orthogonal components, namely the I vector and the Q vector, linked in a complex value of type $I+jQ$, where the I vector is the real component and the Q vector is the imaginary component. Due to their orthogonality, such components can be transmitted together on the wireless channel without interfering. They can also be recovered and assembled at the receiver to reconstruct the original bit-stream. In this context, a typical way to represent complex I-Q signals is through the I-Q plane, as shown in Fig. 1.
@@
In particular, the number of expected I-Q values at the receiver (denoted as $n$) indicates the number of bits that can be recovered through a single complex I-Q value. In general, we can recover $\log_{2}{n}$ bits from $n$ symbols, and thus, with reference to Fig. 1 showing the I-Q plane of a BPSK modulation, we can recover $n=2$ bits, i.e., $\left[i=-1,q=0\right]$ (b=0) and $\left[i=1,q=0\right]$ (b=1). We also notice that the values of the I-Q samples at the receiver always differ from those at the transmitter because of noise introduced by the hardware components of the devices and the fluctuations of the wireless channel between the transmitter and the receiver. To recover the original transmitted symbol, the receiver associates the expected symbol with the received I-Q sample whose distance from the received one is the shortest. Thus, the higher the noise impact, the higher the chance that a received I-Q sample is associated with the wrong expected symbol, leading to an error and, therefore, a higher BER. At the same time, with a given noise profile affecting the communication channel, the higher the modulation order $n$, the higher the amount of expected symbols and, thus, the higher the BER. The rationale described above is adopted by lower-order modulation schemes, e.g., communication links affected by high noise levels, such as satellite transmissions and mobile indoor applications. The intuition driving our work is that the collective displacement of I-Q samples from the expected one can be used to discriminate the presence of various levels of intentional interference, i.e., jamming, affecting the communication link. We will provide more details about our approach in Section V.
III-B Autoencoders
@@
where the symbol $\circ$” represents the composition operator, i.e., $B\circ A(\mathbf{x})=B(A(\mathbf{x}))$, $E$ represents the expectation of the distribution of the input $\mathbf{x}$, ’$A(\mathbf{x})$’ the encoded version of the input, known as the bottleneck of the autoencoder when $p<d$, and finally $\Delta$ is the reconstruction loss function, which measures the distance between the input of the ANN and the reconstruction of the input [42]. For our purposes, in line with many scientific contributions such as [42] and [43], $\Delta$ is the mean-squared-error (mse) function, as defined in Eq. 2 on two reference distributions $\mathbf{x}$ and $\mathbf{y}$.
@@
Traditionally, autoencoders have been used mainly for image generation, particularly for creating sets of images similar to the input ones. However, in the cybersecurity research domain, they are mainly used for anomaly detection. Let $c$ be an autoencoder trained on samples from a probability distribution $P$. Next, let $Q$ be a probability distribution such that $P\neq Q$. Then, we expect $c$ to have a smaller reconstruction error when tested on unseen samples from $P$ than when tested on unseen samples from $Q$. Therefore, the magnitude of the reconstruction error of $c$ in an unseen sample measures the probability that such an unseen sample is not sampled from $P$ but from another distribution [43]. Consequently, we can define a specific error value $\tau$ as a classification boundary. All samples in which the auto-encoder $c$ makes a higher error than $\tau$ can be classified as ’not from $P$’. In the literature, $\tau$ is often referred to as a threshold [44, 45].
@@
In this work, we use autoencoders to build a statistical profile of the channel experienced between the mobile transmitter and the receiver under regular operating conditions. We provide more details in Section V.
IV System and Adversary Model